OpenStudy (anonymous):

4) Solve and check: 5a - 2b = 3 2a = -b

OpenStudy (anonymous):

do you know how to do this type of problem?

OpenStudy (anonymous):

not at all i never learned it

OpenStudy (anonymous):

well if you had an eq with only one variable in it, would you know how to solve it? Such as 3x=15

OpenStudy (anonymous):

yes divide both sides by 3 and get x=5

OpenStudy (anonymous):

right. here you are given in the 2nd eq that -b=2a. so b=-2a. now you can substitute "-2a" in the 1st eq wherever you have a "b"

OpenStudy (anonymous):

5a-2(-2a)3?

OpenStudy (anonymous):

=3

OpenStudy (anonymous):

oh yea i typed it wrong

OpenStudy (anonymous):

but yes you are about to find out what a equals

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

oo i get it :")

OpenStudy (anonymous):

what is 5a-2(-2a) equal to?

OpenStudy (anonymous):

a =3?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

ok im confused..

OpenStudy (anonymous):

-b=6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

a=1/3 so 5*1/3-2b=3 so 5/3-2b=3 so -2b =3-5/3 so -2b= 9/3-5/3=4/3 so b=4/3*(-1/2) sob=-4/6=-2/3

OpenStudy (anonymous):

ok?